解析数論の展望と諸問題

Analytic Number Theory—–Expectations for the 21st Century 研究集会

京都大学数理解析研究所の共同研究事業の一つとして、下記のように研究集会を催しますので、ご案内申し上げます。なお、この研究集会は森田康夫氏代表の科研費からも援助を受けております。

研究代表者 若林 功

（成蹊大 · 工）研究副代表者 村田玲音

（明治学院大 · 経）

記日時： 2000 年 10 月 23 日（月） 13:25 〜 10 月 27 日（金） 17:00 場所：京都大学数理解析研究所４階 420 号室 京都市左京区北白川追分町

プログラム

10 月 23 日（月）

13:25〜13:30 Opening
13:30〜14:30 森田康夫（東北大 · 理）
   不定方程式研究の現状紹介
      An introduction to the theory of Diophantine equations
14:40〜15:40 Yann Bugeaud （Strasbourg 第一大）
      On the Diophantine equation (x^{n} − 1)/(x − 1) = y^{q}
16:00〜17:00 伊藤俊次（津田塾大 · 学芸）
      三次 Pisot 数による β 展開と Thurston-秋山タイリング
   Cubic Pisot β-expansions and Thurston-Akiyama Tilings
17:10〜17:40 桂田昌紀（慶応大 · 経）
      Hurwitz ゼータ関数の導関数の特殊値
      Specific values of derivatives of Hurwitz zeta-functions

10 月 24 日（火）

9:30〜10:30 松本耕二（名大 · 多元数理）
      約数問題と Riemann ゼータ関数の二乗平均値問題についてのサーヴェイ
      A survey on the divisor problem and the problem on the mean square
      of the Riemann zeta-function
10:45〜11:45 P.D.T.A. Elliott （Colorado 大）
      Probabilistic number theory and its ramifications
13:30〜14:30 荒川恒男（立教大 · 理）
      SL(2, R) の Selberg 跡公式と低ウェイト尖点形式空間の次元などについて
      Selberg trace formula for SL(2, R) and dimensions of certain spaces
      of cusp forms with small weights
14:40〜15:30 二宮祥一（東工大 · 理財工学研究センター）
      金融の現場における一様分布論の応用について

On applications of uniformly distributed sequences to finance problems

15:50〜16:40 見正秀彦 (名大 · 多元数理)
      量指標を持つ L 関数の Universality theorem
      The universality theorem for L-functions associated with
      Grössencharacters
16:50〜17:40 田村純一
      或る高次元連分数の p 進値
      p-adic values of certain continued fraction of high dimension

10 月 25 日（水）

9:30〜10:30 川田浩一（岩手大 · 教育）
      二十世紀のワーリング問題とその周辺
      Waring’s problem and variants in the twentieth century
10:40〜12:10 本橋洋一（日大 · 理工）
      素数分布 — 振り返って
      Distribution of Prime Numbers — looking back a bit personally
13:30〜14:30 Peter Bundschuh （Köln 大）
      Arithmetical properties of functions satisfying linear q-difference
      equations: a survey
14:40〜15:10 田中孝明（慶応大 · 理工）
      ある種の空隙的ベキ級数の値の超越性
      Transcendence of the values of certain lacunary series

15:20〜16:10 古屋淳（名大 · 多元数理）
      二次体のイデアルの個数関数を含む指数和について
      On exponential sums involving the ideal counting function
      in quadratic fields
16:20〜17:10 竹内良平（都立大 · 理）
      二項方程式を満たす原始根の分布について
      On the distribution of primitive roots satisfying a binomial equation

レセプション 18:00〜 於 京大会館 Reception at Kyoto University House 18:00〜

10 月 26 日（木）

9:30〜10:30 平田典子（日大 · 理工）
     楕円曲線上の対数一次形式についての最近の進展と問題
      Linear Forms in Logarithms on Elliptic curves:

a survey on recent results and open problems

10:45〜11:45 Attila Pethő （Debrecen 大）
      The first 100 years of algorithmic theory of diophantine equations
13:30〜14:30 岡崎龍太郎（同志社大 · 工）
      CM-体の類数１問題の進展について
      On progress in class number one problem of CM-fields
14:40〜15:40 Marc Huttner （Lille 第一大）
      On a paper of Hermite and Diophantine approximation

of Abelian integrals

16:00〜16:50 名越弘文（慶応大 · 理工）
      保型 L-関数の族の universality theorem
      The universality of the family of automorphic L-functions
17:00〜17:30 鹿野健（山形大 · 教育）
      Positivity of certain trigonometric sums

10 月 27 日（金）

9:30〜10:30 江上繁樹（富山大 · 工）
      Dedekind 和の一般化
      Generalizations of the Dedekind sum
10:45〜11:45 小山信也（慶応大 · 理工）
      ３次元 Eisenstein 級数の量子エルゴード性
      Quantum elgodicity of Eisenstein series for arithmetic 3-manifolds
13:30〜14:30 秋山茂樹（新潟大 · 理）
      数論的アルゴリズムに付随するタイル張り
      Tilings attached to the arithmetic algorithm
14:40〜15:10 吉元昌己（京大 · 数理研）
      ゼータ関数値のラマヌジャン急収束級数
      On Ramanujan’s rapidly convergent series for Riemann zeta-values
15:20〜16:00 市原由美子（名大 · 多元数理）
      指標付き Rankin-Selberg L 関数の係数に関する Riesz 和について
      On Riesz mean for the coefficients of the twisted Rankin-Selberg
      L-functions
16:10〜16:50 知念宏司（大阪府立大 · 総合科学），村田玲音（明治学院大 · 経）
      a (mod p) の剰余位数の分布について
      On a distribution property of the residual order of a (mod p)